OpenStudy (anonymous):

Can someone please help me with this problem? Juana was training to run a marathon. During the three day period before the race she decided that she would train for a total of 11 hours/ If she trained for 2 3/5 hours on the first day and 2 9/10 hours on the second day, how many hours would she need on the third day?

OpenStudy (anonymous):

i believe juana wound need to run 2 3/4 on the third day

OpenStudy (sleepyjess):

2 3/5 = 2 6/10 2 6/10 +2 9/10 = 5 5/10 or 5 1/2 11 (total hours she wants to train) - 5 1/2 (hours already trained) = ? can you solve the last part? @cludicate1223

OpenStudy (anonymous):

@Cosmichaotic